OpenStudy (anonymous):

4. Find the missing value to the nearest hundredth. sin_7/12 plz help me

OpenStudy (anonymous):

so you need to just turn it into a decimal?

OpenStudy (anonymous):

yeah if u need the answers i can give them to u

OpenStudy (anonymous):

oh no i was just making sure i understood the question thanks though

OpenStudy (anonymous):

kk ty for helping me i am no good with math lol

OpenStudy (anonymous):

are you allowed to use a calculator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so if you want 7/12into a fraction just divide 7 into 12 then you have a decimal

OpenStudy (anonymous):

35.69° 30.26° 80.16° 54.31° thoes are the choices though

OpenStudy (anonymous):

just plug '\(\sin(\frac{7}{12})\)' into your calculator and reduce the decimals to 2.

OpenStudy (anonymous):

it gives me 0.010180911

OpenStudy (anonymous):

multiply that by 360 to get the degrees.

OpenStudy (anonymous):

i wonder why the choices are in degrees....are you sure the question is written entirely and correctly?

OpenStudy (anonymous):

@cthomasknight, why would you do that?

OpenStudy (anonymous):

yes i coppied it from my test itself so ik it is right

OpenStudy (anonymous):

Is the question written right?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

it says find the missing vaule to the nearst hundredth sin_____=7/12

OpenStudy (anonymous):

there you go. That's what we thought.

OpenStudy (anonymous):

that makes much more sense. Do you know how to do it?

OpenStudy (anonymous):

no i suck at math

OpenStudy (anonymous):

all you gotta do is put '\(\large \sin^{-1}(\frac{7}{12})\)' into your calc.

OpenStudy (anonymous):

oh ok lol

OpenStudy (anonymous):

so i would do the same thing with the cos______=9/20

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ty lol

OpenStudy (anonymous):

math is easy if you know what you're looking for ;) good luck

OpenStudy (anonymous):

i dont lol

OpenStudy (anonymous):

haha rome wasn't built in a day. Keep trying and you'll start to like it, hopefully :)

OpenStudy (anonymous):

i wish i could choose both of u as the best response lol

OpenStudy (anonymous):

it's the thought that counts. Thanks :)

OpenStudy (anonymous):

np yall were really nice to help me lol

OpenStudy (anonymous):

Good luck with the rest of the test :)